OpenStudy (loser66):

Prove: \(\dfrac{cos2A}{2}+1=\dfrac{-cosA}{2}+\dfrac{1}{4}\) Please, help

OpenStudy (rational):

Hint cos(2A) = cos(A + A) = cosAcosA - sinAsinA = cos^2A - sin^2A = 2cos^2A-1

OpenStudy (loser66):

Can you give me more. I am sorry for being blank right now.

OpenStudy (rational):

I don't have pen and paper handy... i would just mess with that angle sum identity but ur equation/identity doesn't look correct hmm

OpenStudy (loser66):

Another question relates on it: Let y = 2pi/5 2y = 4pi/5 \(cos 2y = 2 cos^2 y -1\) then \(cos^2y -\dfrac{1}{2} cos2y-\dfrac{1}{2}=0\) How to write it w.r.t y only. I know the answer is \(cos^2y+1/2 cos y -1/4 =0\) but don't know how to do

OpenStudy (loser66):

@tkhunny